OpenStudy (anonymous):

I need help applying chain rule ! Suppose that f(x)=4x^2(3-5x)^3 Find the equation for the line to the graph of f at x=-2 I thought of using (f'g-fg')/ g^2

OpenStudy (nincompoop):

? how is that a chain rule?

OpenStudy (nincompoop):

chain rule would be the derivative of the outer function multiplied by the derivative of the inner function

OpenStudy (nincompoop):

so we have h(x) = (x+2)^3 h' = 3(x+2)^2 * 1 here's the technique (x+2) = u; so u^3 = (x+2)^3 which leads us to the derivative of u^3 to be 3u^2 = 3(x+2)^2 then we still need to obtain the derivative of the inner function (x+2)' = 1

OpenStudy (anonymous):

Where did you get (x+2) ?

OpenStudy (nincompoop):

just my own example so I won't end up doing your homework for you. I teach concepts not give out answers

OpenStudy (anonymous):

Sorry I noticed, I just wanted to understand.

OpenStudy (anonymous):

Thank you for your example I'll work out the problem from here and see if I can come up with the correct answer :)

OpenStudy (nincompoop):

okay let me know if you need further clarification.

OpenStudy (anonymous):

I understand the example but I'm having trouble with the 4x^2

OpenStudy (nincompoop):

you mentioned I thought of using (f'g-fg')/ g^2 this is quotient rule btw

OpenStudy (anonymous):

In my case f(x)=4x^2(3-5x)^3 f(x)=4x^2 ----> f'(x)=8x g(x)=(3-5x)^3 --> g'(x)=3(3-5x)^2 and I would get g'(f(x))f'(x) so 3(8x)^2*(8(3(3-5x)^2) I really am trying, I've been stuck on this problem for a bit now :/

OpenStudy (nincompoop):

so after you do the chain rule, 4x^2 juts indicates that you multiply it leading to a product rule

OpenStudy (anonymous):

Is my work right so far ?

OpenStudy (nincompoop):

your chain rule is incomplete (3-5x) needs a derivative as well

OpenStudy (anonymous):

the derivative of (3-5x) would be -5

OpenStudy (nincompoop):

k so make sure you put them together first

OpenStudy (nincompoop):

3(3-5x)^2 * -5

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

so I would have -15(3-5x)^2

OpenStudy (nincompoop):

go on

OpenStudy (anonymous):

3(8x)^2 *(8(-15(3-5x)^2) ?

OpenStudy (nincompoop):

where did you get the constant 8?

OpenStudy (anonymous):

I mean to write 8x which is the derivative of 4x^2

OpenStudy (nincompoop):

rewrite it clearer

OpenStudy (anonymous):

so my work in the top was not right ? I took the derivative of 4x^2 = 8x the derivative of (3-5x)^3 = 3(3-5x)^2 and you told me to take another derivative of (3-5x) = -5 I've understood this but now I'm getting lost into how to put all this together

OpenStudy (nincompoop):

product rule: u'v + uv' f(x)=4x^2(3-5x)^3 u = 4x^2; u' = 8x v = (3-5x)^3; v' = 3(3-5x)^2 * -5 or -15(3-5x)^2 now just do the product rule

OpenStudy (anonymous):

8x(3-5x)^3 + (4x^2)(-15(3-5x)^2) would be my setup ? For this particular problem to find my slope for the tangent line equation that I'm supposed to find can I automatically plug in x=-2 right away ?

OpenStudy (nincompoop):

correct set up you can plug x = -2 right away if you want, but simplifying would be the best way to go

OpenStudy (anonymous):

Oh okay thank you, thank you so much for your time and help !! I understand this concept a bit better :)

OpenStudy (nincompoop):

be sure to watch this http://youtu.be/6kScLENCXLg?t=2m36s and this one https://www.youtube.com/watch?v=zwuQAOFvIV4

OpenStudy (anonymous):

Okay :) thank you .

OpenStudy (nincompoop):

they are really good examples